Shift summary for owl 8/22/99

Jason Hamm (jhamm@physics.arizona.edu)
Sun, 22 Aug 1999 04:59:02 -0700 (MST)


Shift Summary for owl shift, 08/22/99

On Shift: J. Hamm, T. Alexopolous

Hours of beam available at all intensities: 7
Hours of beam available at nominal intensity: 7
Hours of E832 data (special runs not included): 6.75
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: Changing tapes, stream 5 errors
========================================================================
Runs ended this shift:

14298 - 23.7 K vac 2pi0

Tapes brought to FCC this shift: QK(S,C,K)303-305, QKB106
QKA015, QKSM13, QKCC14,
QKNL24, QKN026, QKCR06, QKPE12
========================================================================
Accesses during shift: none
========================================================================
Overview of shift: Incredibly smooth running at ~6.8 E12
========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14298 0:10 Online histos show that 2V1 source timing has
shifted from -6.9 ns to -7.2 ns, resulting in an
unhappy face. RSK is consulted. See comments
below.

14298 1:55 Stream 5 errors, miss a spill.

3:25 Tapes fill, run 14298 ends. It appears that the
histograms were not saved, even though the run
stopped gracefully. Or are they in a new directory?

3:37 Start Beam_832 run 14299. No ktevana errors at
startup!

3:40 Ktev3c is 91% full. Page Sydney, who has us
delete the caches for all the onsplit tapes listed
above. Crisis averted.

14298 7 3:45 Stream 5 errors, miss this spill

4:15 Onsplit requests QKA016 on ktev3

14298 64 5:00 Laser crate DYC full, miss this spill

14298 94 5:45 Crate 53 DYC busy, miss this spill
========================================================================
Follow up Notes on Previous Shift
========================================================================
Items to Follow up on Next Shift

rf shift is ok for now, but a larger shift could start causing hcc
inefficiencies. If inefficiencies start to occur, check the 2V1 source
timing plot (in trg#summary) to see if the timimg has shifted further.

Also, a few times the da was dead for 1-5 seconds at the beginning of a
spill, resulting in a daqdead error. We were not able to find any
possible cause for this.